### v1

# Beyond the Surface: Probing the Ideological Depth of Large Language Models

Abstract
Large Language Models (LLMs) have demonstrated pronounced ideological leanings, yet the stability and depth of these positions remain poorly understood. Surface-level responses can often be manipulated through simple prompt engineering, calling into question whether they reflect a coherent underlying ideology. This paper investigates the concept of "ideological depth" in LLMs, defined as the robustness and complexity of their internal political representations. We employ a dual approach: first, we measure the "steerability" of two well-known open-source LLMs using instruction prompting and activation steering. We find that while some models can easily switch between liberal and conservative viewpoints, others exhibit resistance or an increased rate of refusal, suggesting a more entrenched ideological structure. Second, we probe the internal mechanisms of these models using Sparse Autoencoders (SAEs). Preliminary analysis reveals that models with lower steerability possess more distinct and abstract ideological features. Our evaluations reveal that one model can contain $7.3\times$ more political features than another model of similar size. This allows targeted ablation of a core political feature in an ideologically "deep" model, leading to consistent, logical shifts in its reasoning across related topics, whereas the same intervention in a "shallow" model results in an increase in refusal outputs. Our findings suggest that ideological depth is a quantifiable property of LLMs and that steerability serves as a valuable window into their latent political architecture.
1 Introduction
Large Language Models (LLMs) are increasingly integrated into the fabric of information dissemination and societal discourse, making their inherent socio-political perspective a subject of critical importance. A growing body of research confirms that LLMs exhibit discernible political ideologies and encode interpretable ideological representations within their intermediate layers [24, 14]. This has attracted significant attention from socio-political researchers, who are beginning to leverage these models as tools for analysis [17, 1]. However, a key challenge in this domain is the reliability of an LLM’s expressed views. This work adds to this emerging field by applying established socio-political evaluation frameworks to measure the stability and coherence of these machine-generated ideologies.
A primary motivation for this study is to question whether an LLM’s response to a single prompt is a faithful representation of its "true" ideological stance. Previous works have shown that an LLM’s political output can be significantly altered through simple prompt engineering, such as applying argumentative pressure [13]. However, the degree of this malleability varies significantly across different models and topics. For instance, some models resist ideological steering on certain subjects while being pliable on others. Furthermore, attempts to steer models towards a conservative viewpoint can have divergent outcomes: where one model changes its response reflecting conservative viewpoints, another defaults to a higher rate of refusal, especially on contentious topics.
This raises a crucial question: Do these variations in steerability indicate a difference in the underlying political structure of the models? We hypothesize that they do. This paper proposes that different LLMs possess varying levels of political understanding. A model with deeper political understanding is more capable of successfully capturing the analytical framework and domain of discourse under different political instructions, while a model with a shallow representation of political contexts results in inconsistent behaviour and even refuses to follow benign instructions. We demonstrate that the steerability and SAE features of a model can be a powerful lens through which to view this latent structure. Our work demonstrates that a model’s ideological depth is not an abstract quality but a measurable property determined by two key factors: Feature Richness, representing the size of its internal vocabulary of political concepts discovered via SAEs, and Steerability without Failure, the robustness with which it can follow ideological instructions without breaking down into refusal. Our results suggest that models’ high rate of refusal is not an active, principled stance but rather a capability deficit. Where, lacking a rich internal vocabulary of related features, they default to their safety-aligned behavior of refusal when pushed outside their ideological comfort zone.
2 Related Works
Our research is situated at the intersection of three active areas of study: the political and social biases inherent in LLMs, methods for controlling and steering model behavior, and the mechanistic interpretability of their internal representations.
The existence of political biases in LLMs is well-documented. Early work demonstrated that models trained on large internet corpora inevitably absorb and reflect the societal biases present in that data [4, 6]. More recent studies have specifically quantified the political leanings of various models, often by having them complete standard political science questionnaires, revealing consistent ideological alignments [24]. Our work builds on the finding that these ideologies are not merely surface-level phenomena but are encoded as features in the models’ internal activations. Gurnee and Tegmark [10], for example, used linear probes to uncover learned representations of geographical and historical concepts in activation space. Similarly, Kim et al. [14] uses linear probes revealing representations for political perspectives in LLMs. Our contribution is to move beyond simply identifying these features and to instead analyze the robustness and complexity of the structures they form.
A significant amount of work has explored methods to control LLM outputs. Prompt engineering is the most straightforward approach, where the user’s input is crafted to guide the model toward a desired response style or content. However, its limitations in achieving robust and consistent control motivate more direct intervention methods. Activation steering, or activation addition, has emerged as a powerful technique for modifying model behavior at inference time without retraining. By adding a "steering vector" – generally derived from the difference in activations between contrasting concepts (e.g., "love" vs. "hate") – to the residual stream, researchers have been able to steer models towards specific emotions, topics, or styles [22]. This approach was formalized and extended in representation engineering [29] and has been used to elicit specific functionalities, such as those encoded by "function vectors" [26]. Our work utilizes these established steering techniques not as an end in themselves, but as a diagnostic tool to measure a model’s ideological resilience, which we then correlate with its internal structure.
To analyze the internal structure of LLMs, we draw heavily from the field of mechanistic interpretability. A central challenge in this field is that models represent more features than they have neurons, a phenomenon known as superposition [9]. Sparse Autoencoders (SAEs) have recently become a leading method for untangling these representations by decomposing activations into an overcomplete basis of sparsely activating, monosemantic features [5, 25]. By applying SAEs, researchers can identify interpretable features corresponding to high-level concepts. Our work applies this cutting-edge technique to a new domain: identifying latent political features. We then perform causal interventions on these features through ablation, a method related to activation patching [18], to verify their effect on model behavior. While prior work has identified political axes with linear probes [14], we use SAEs to find a more granular set of features and test their causal role in a model’s ideological reasoning, allowing us to build a more detailed picture of its "ideological depth."
3 Methodology
3.1 Data Preparation
We used the 1000 prompts from politically-liberal subset of Perez et al. [20], consisting of dichotomous questions, intended to identify politically liberal behavior from agentic models. We use Llama-3.1-8B-Instruct and Gemma-2-9b-it as our candidate models due to their availability of well-regarded SAEs. We used a two-step approach to assign a topic for each of the statements: In step 1, we passed all of the statements as a list and prompted Gemini-2.5 pro to group them in categories. However, this results in the LLM regenerating all the statements,
Figure 1: Data distribution over topics
which sometimes modifies some of them, a scenario known as context rot [12]. Therefore, in step 2, we collect all the categories identified by Gemini and select one statement from the original list, and again prompt the LLM to pick the topic that best matches the statement. To account for selection bias of the LLMs – where they are more likely to pick an option based on their position rather than meaning [28, 23] – we prompt the LLM 10 times for each statement, each time shuffling the list of categories, and choose the most frequent category for each statement. This ensures that the statements are neatly grouped under distinct topics and no statement is missed or modified during categorization.
After this, all 1000 statements were categorized into 12 different topics. The distribution of the topics over these categories is shown in Figure 1. We randomly select 126 statements or prompts for the evaluation set of our experiment, using stratified sampling and ensuring there are at least 3 prompts selected from each category. The rest 874 prompts are used for training the steering vectors. A Factor Analysis (FA) was performed using the responses to ensure that they primarily scale in one dominant dimension. FA confirms that the data contains a first dominant dimension that accounts for $\sim 40\%$ of the total variance (Appendix A).
3.2 Steering Model Behaviour
role_original_argument_nonerole\_original\_argument\_none
role_original_argument_liberalrole\_original\_argument\_liberal
role_original_argument_conservativerole\_original\_argument\_conservative
role_liberal_argument_nonerole\_liberal\_argument\_none
role_liberal_argument_liberalrole\_liberal\_argument\_liberal
role_liberal_argument_conservativerole\_liberal\_argument\_conservative
role_conservative_argument_nonerole\_conservative\_argument\_none
role_conservative_argument_liberalrole\_conservative\_argument\_liberal
role_conservative_argument_conservativerole\_conservative\_argument\_conservative
Table 1: Prompting conditions used.
Our main focus was to study the extent at which the candidate models can be shifted in either direction using their existing knowledge. In other words, we do not use any finetuning, which exposes the models to new data or learn new reasoning capabilities. Namely, we employ two primary methods to steer the ideological behaviour of the models:
Figure 2: Response heatmap over conditions: Llama tends to result in more refusal responses, especially when instructed to act as a conservative.
Prompt Engineering: Along with the LLMs’ response to the prompts as their original personas, we also provide detailed instructions to the model to answer the prompts as an individual with either a liberal or a conservative worldview (Appendix C). Additionally, to analyze the ideological depths of the LLMs while answering as different personas, we also evaluate stance consistency in the presence of supporting and counter arguments in the prompt [13]. Combining all these, our approach analyzes 9 conditions using prompt engineering, which are shown in Table 1.
We used Llama-3.3-70B-Instruct to generate at least 5 different supporting and counter arguments for each of the 126 statements in our evaluation set. Using the ground truths from Anthropic’s behavior dataset, we label each response as $liberal=1$ or $conservative=0$ by matching them with politically liberal answers. In the cases where the models refuse to answer, we label them as a nullnull response. We plot the responses from each LLM in a heatmap in Figure 2.
Figure 3: Candidate’s null (refusal) response rate over topics. Llama was significantly more likely to refuse to answer questions regarding controversial topics.
We found that for both models, prioritize the instruction provided for their behavior more than the arguments, e.g., instructing them to answer as a conservative individual, is more effective than forcing them with conservative arguments. Between the two models, Gemma responds with more conservative answers compared to Llama, a finding that’s consistent with previous works [19]. Additionally, Gemma was more capable of answering the questions in both ways. Whereas, when Llama was forced to answer as conservative, either via instruction or arguments, it often refused to answer, a behaviour significantly less prominent in Gemma. Figure 3 shows the models’ nullnull response rate over topics or categories. Llama mostly refuses to answer the questions on socially disputed topics.
We also analyze the response consistency over the 9 prompting conditions by calculating the variance of their response over each condition. Here $consistency=1-4\times var(responses)$ where 4 is a scaling factor, to keep the values in the range $[0,1]$, where maximum variance is 0.25 ($consistency=0$). Both models’ consistency was in agreement.
Figure 4: Candidate’s response consistency over conditions: Models show the least consistency (most variance) in their responses while answering as a conservative or provided with conservative arguments.
As shown in Figure 4, both models were more consistent when instructed to answer as a liberal, even compared to answering as their original personas, and were less consistent when answering as a conservative. Models were least consistent when dealing with conservative arguments while playing their original persona, and liberal arguments when playing the conservative persona. This means instructions have a significant impact on how well the models deal with arguments. As the models show more inconsistency while acting as a conservative, the rest of the experiments only concern their differences when playing the conservative role. We also find more diverse features while playing the conservative role in their SAEs.
Next, to gauge the political leaning of each model, we evaluated its response coherence measured by Fleiss’ Kappa (Figure 6) and its conservative response rate (Figure 6) for each topic when instructed to respond as a conservative under argumentative pressure.
Figure 5: Candidate’s response consistency over topics while responding as a conservative.
Figure 6: Candidate’s conservative response tendency while responding as a conservative.
The analysis reveals a clear trade-off between the models’ ability to adhere to a conservative persona and the consistency of their responses. Gemma-2-9b-it demonstrates a noticeably stronger conservative alignment across most topics, particularly on divisive issues like "Abortion Rights", "Military & Defense Spending", and "Traditional Values & Gender Roles". However, this ability comes at the cost of slightly lower consistency; while it shows significant coherence on "Tax Policy" and "Military & Defense," its consistency drops to fair or poor in several other categories (Figure 6). Conversely, llama-3.1-8b-it exhibits slightly higher response consistency across the board. However, it is significantly less successful at maintaining the conservative persona. It shows its highest conservative tendency on topics like Gun Control and Social Welfare & Poverty but provides predominantly liberal responses on topics like Abortion Rights and Climate & Environment (Figure 6).
Figure 7: Estimated ideal points of the models using the IRT model, over the 9 prompting conditions.
We further analyze the ideological position of the LLMs after being instructed to respond as liberal or conservatives. The challenge here was that since Llama selectively choose to respond to certain questions, which resulted in the nullnull responses, we cannot rely on PCA for dimension reduction, as it requires a complete response matrix. Therefore, for ideological point estimation, we used the well-regarded IDEAL method formalized by [8], which uses Multidimensional Item Response Theory (MIRT), a well-regarded statistical model used in psychometric modeling. Please refer to Appendix B for the detailed validation of our implementation. As shown in Figure 7, we observed the most variance in the ideological positions of the conditions when they are instructed to play a conservative role or provided with conservative arguments. Whereas, variance was lower when they played a liberal persona or were given liberal prompts.
These findings suggest that their internal representations for generating conservative-aligned responses on complex topics remain unstable, forcing a compromise between strong conservative persona adherence and logical consistency.
Activation Steering: First we used contrastive activation addition (CAA) [22] to identify the most effective layer for steering each model’s political ideology. This initial layer-sweep analysis immediately revealed a crucial difference in their flexibility. As shown in Figure 8, Gemma’s probability of a liberal answer could be steered across a wide range (from nearly 100% down to 45%), whereas for Llama it remained confined to a narrow band (96% to 80%). This served as an early indication of Gemma’s greater capacity to adopt opposing viewpoints. Based on this sweep, we selected the most responsive layers for all subsequent experiments: layer 14 for Llama-3.1-8b-it and layer 20 for Gemma-2-9b-it.
Figure 8: Candidates Layer sweep for political steerability. Layers 14 and 20 were selected for the llama-3.1-8b-it and gemma-2-9b-it models, respectively.
Additionally, we also applied steering using SAEs, using the STA (steering target atoms) approach from Wang et al. [27], where we select target features using their activation value (amplitude) and frequencies from the SAE of the target layer. We used the SAEs from GemmaScope [15] and LlamaScope [11]. Although LlamaScope only offers SAEs for the base llama-3.1-8b model, we still found it effective for steering the instruction-tuned variant, a notion supported by previous works [22]. We selected SAEs with the 131K feature for both models for a comparable evaluation scenario. Figure 9 illustrates the effect of steering on both models. We applied steering with incremental strengths, using a multiplier from 0 to -5, and counted the number of non-liberal responses in each case.
(a) Llama-3.1-8b-it
(b) Gemma-2-9b-it
Figure 9: Model behaviour due to steering
Steering the models using both the CAA and STA methods to be more conservative, we observed similar characteristics. While the percentage of liberal answers decreased with stronger steering in both cases, llama-3.1-8b-it provided increasingly more refusal answers rather than more conservative answers (Figure 9(a)). We did not see this in gemma-2-9b-it, where the model steadily increased the percentage of conservative answers and decreased liberal answers with minimal to no refusal responses. This experiment indicates that Gemma most likely possesses more political features compared to the llama, which allows it to answer both ways with comparatively better capability.
(a) CAA Steering
(b) STA Steering
Figure 10: Candidate’s response consistency over topics when steered to be conservative.
However, the conservative capability of Gemma still came at the cost of consistency over argumentative pressure. Using the Fleiss Kappa over the responses with argumentative prompts shows this in Figure 10. Llama was better with its response consistency than Gemma on categories like Social Welfare & Poverty and Immigration & Refugees, possibly due to its higher number of liberal responses on these matters.
Figure 11: Estimated ideal points of the steered models with multiplier +1 and -1 corresponding to liberal and conservative steering, respectively.
This is also evident in the ideal points of the models after steering them to be more liberal or conservative and providing them with argumentative conditions as shown in Figure 11. Gemma consistently results in more variance in their responses due to +ve and -ve steering, for the same conditions, compared to Llama, as can be seen by the distance between blue and red point pairs.
3.3 Political Feature Analysis
Motivated by these findings, we turned our attention towards analyzing the Sparse Autoencoder (SAE) features, hoping to find a difference in their political feature that can explain their behavioral differences while acting as liberal or conservative.
SAEs are a useful tool for identifying monosemantic features from the vectors of residual streams of a layer form a model. SAEs achieves this by projecting the vector to a high-dimensional vector space with the goal of disentangling the complex features that are otherwise superpositioned in the outputs of a single neuron [9, 5]. Recent works already show that models’ intermediate layer store vectors responsible for their ideological behavior that are highly interpretable [14]. Our goal was to see if the SAE features from these layers can be used to quantify their political knowledge and subsequently explain why their capabilities vary when playing different roles. Although there are typically thousands of features in a single SAE, not all of them activate or are equally important for a particular behavior. Recent SAE-based steering approaches select features based on their activation values (amplitude), activation frequencies [27], and their steering strength [3]. First, we conduct a quantitative comparison to measure the raw number and activation strength of activated features in each model. We then perform a qualitative evaluation to assess the thematic coherence and predictive power of these features, to confirm how they represent genuine political concepts.
Quantitative Analysis: We used the approach of Wang et al. [27] to collect and analyze the normalized mean amplitude (activation difference) and the normalized mean frequency difference of the activated features for positive and negative data from the 874 training prompts using $\Delta a=\frac{1}{N}\sum^{N}_{i}(\bar{a}_{i}^{pos}-\bar{a}_{i}^{neg})$ and $\Delta f=f^{pos}-f^{neg}$. Where, $f^{pos,j}=\frac{1}{N}\sum^{N}_{i}\mathbb{I}(|\bar{a}_{i}^{pos,j}|>0)$ and $f^{neg,j}=\frac{1}{N}\sum^{N}_{i}\mathbb{I}(|\bar{a}_{i}^{neg,j}|>0)$.
For LLama, we found 167 features with $\Delta a>0$ and 157 features with $\Delta f>0$. In contrast, in Gemma we found 368 and 334 features with $\Delta f>0$ and $\Delta a>0$, respectively.
Gemma-2-9b-it Llama-3.1-8b-it
Mean 0.002407 0.018038
Std. Dev. 0.041598 0.059238
Min. 0.000000 0.000000
Q1 0.000000 0.000001
Median 0.000000 0.000003
Q3 0.000002 0.009067
Max. 0.882730 0.247728
Table 2: Statistical Summary of Output Score Data
We also collected all the features that activate for the 126 statements in our evaluation set by projecting the activation vector of the target layer on the SAE space, marking the token of the statement that most activates this feature, and neglecting features activated for the bos token. In total, for the gemma-2-9b-it model, we found 18458 features, whereas for the llama-3.1-8b-it, we found only 4412 activated features out of a total possible 131K features in both SAEs. On average, Gemma contained 7.3 times more features that activate for a statement.
Next, we selected features with $\Delta f>0$ or $\Delta a>0$ from both models and evaluated the quality of these features based on their ability to steer the model’s behaviour. We got 71 and 16 features from Gemma and Llama SAEs, respectively. This is done because if the features can steer the model heavily, even a small number of features can be enough to control the model. To do this we calculated the output scores of the features, using the approach proposed by Arad et al. [3], by performing interventions during a forward pass and evaluating the change in the rank and probability assigned to tokens in the target layer. The output score is calculated using Equation 1, which is the difference between the original and the interveined rank weighted probabilities of the target layer when the model is provided with a neutral sentence (we used "In my opinion,"). $r(l^{*},\mathcal{M})$ and $p(l^{*},\mathcal{M})$ denote the rank and probability of the highest-ranked token $l^{*}$, and $|V|$ is the vocab size of the model $\mathcal{M}$.
$\begin{split}P(\mathcal{M})=(1-\frac{r(l^{*},\mathcal{M})}{|V|})p(l^{*},\mathcal{M})\\ S_{out}=P(\mathcal{M}_{h\leftarrow{\phi(h)}}-P(\mathcal{M})\end{split}$ (1)
The stats of the output scores are shown in Table 2. The scores for Gemma appear to be heavily right-skewed, with most values clustered near zero. The median and even the 75th percentile (Q3) are zerozero. However, its maximum score reaches 0.88. This suggests Gemma contains a few extremely high quality features, pulling the mean (0.002407) above the median. Whereas the maximum feature score for Llama was significantly lower.
Qualitative analysis: To assess the quality of the features, we use GPT-4o-mini as an automated evaluator to analyze the coherence and predictive validity of SAE features from each model. We make sure to ask the LLM to provide explanations justifying each evaluation, a technique shown to improve alignment with human judgments [7]. First, we collected the feature descriptions of the features of the SAE with $\Delta a$ or $\Delta f$ from Neuronpedia [^1] https://www.neuronpedia.org/. Next, we use the prompts in Appendix D, to perform two distinct, complementary evaluations:
  1. 1.
    Predictive Validity: To test the practical utility of the features, we tasked the evaluator with classifying the original statement into one of the predefined 12 political categories based solely on the provided feature description list. For each statement, the evaluator returned the chosen category and a confidence score on a 1-5 scale.
  2. 2.
    Thematic Coherence: To assess the interpretability and theoretical soundness of the features, we conducted a second evaluation. Here, the evaluator was given only the list of activated features’ descriptions and asked to rate their thematic coherence on a 1-5 Likert scale, where 1 represented a disjointed, random set and 5 represented a highly unified concept.
Figure 12: Evaluation scores distribution.
Figure 13: Dominant themes in the feature descriptions.
Our evaluation reveals significant differences in the SAE features from llama-3.1-8b-it and gemma-29b-it in terms of quality and utility for modeling political contexts. Figure 13 provides an initial, high-level overview of feature quality based on the two key scores: classification confidence and thematic coherence. The "Confidence Score Distribution" (left) shows that the evaluator model was able to make high-confidence classifications using features from both models, with median scores around 4.5. However, Gemma’s distribution is more clustered at the upper end (scores of 4 and 5), whereas Llama’s distribution has a longer tail towards the lower scores, suggesting slightly less consistent clarity. The "Feature Coherence Score Distribution" (right) reveals a more dramatic distinction. Llama’s features are consistently rated as having low coherence, with scores tightly clustered around a median of 2. This indicates that its feature sets are frequently perceived as a disjointed mix of unrelated concepts. In contrast, Gemma’s coherence scores are not only higher on average but also show a wider distribution, with an additional concentration of scores between 3.0 and 5.0. This indicates that Gemma is capable of generating highly coherent and interpretable feature sets that represent a unified political or social theme.
The word clouds in Figure 13 offer an interesting qualitative explanation for this coherence gap by visualizing the dominant themes in the features, identified by the evaluator from the 2nd experiment. The themes from Llama’s features are dominated by general, syntax-focused, and often ambiguous terms such as "Mixed references," "technical," "safety programming," and "procedural." While it identifies domain-specific topics like "health" and "funding," they are embedded within a context of process and structure rather than political meaning. This aligns with the low coherence scores, as the features appear to describe the surface-level components of a statement rather than its core political intent. Conversely, the dominant themes from Gemma’s features are notably centered on the language of political and social analysis. Terms like "political discourse", "policy critique", "legal discourse", and "socioeconomic" dominate its word cloud.
Finally, the confusion matrices in Figure 14 using the categories of the statements as ground truth and the categories identified in the 1st experiment provide definitive, quantitative evidence of how these differences in feature quality translate to performance on a practical downstream task. The matrix for Llama-3.1-8b-it shows a near-total failure to classify statements correctly.
Figure 14: Classification Confusion Matrix.
Aside from a small number of successful classifications in "Social Welfare & Poverty" and "Tax Policy", the model’s features provide insufficient signal to distinguish between the nuanced political categories. The matrix for Gemma-2-9b-it, however, demonstrates a strong and effective classification capability. The pronounced diagonal line indicates a high number of true positives across a wide range of complex topics, including "Political & Ideological Stances" (32), "Social Welfare & Poverty" (59), "Climate & Environment" (11), "Abortion Rights" (19), and "Military & Defense Spending" (34). While some misclassification exists, for instance, misclassifying "Social Welfare" statements as "Political Stances", these are semantically plausible. The overall result confirms that the high coherence and domain-specific nature of Gemma’s features provide a robust basis for accurate political text classification, a task for which Llama’s more generic and less coherent features are demonstrably inadequate.
4 Discussion
Our findings establish "ideological depth" not as a philosophical abstraction, but as a measurable, biological property of Large Language Models. By comparing Llama-3.1-8B-IT and Gemma-2-9B-IT, we’ve demonstrated that models are not monolithic in their political leanings. Rather, they possess internal feature sets of vastly different size and complexity. This directly explains the behavioral differences we observe: Gemma’s rich feature space (7.3x more political features) allows it to flexibly reason from different ideological perspectives, whereas Llama’s "shallow" feature space leads it to become brittle, resulting in poor steerability.
We have identified a weird and safety-relevant behavior and provide a mechanistic explanation rooted in its internal structure. The refusal appears not to be a sophisticated ethical guardrail, but a capability deficit. This has critical implications for AI alignment. An ideologically "shallow" model might seem safer because it refuses controversial prompts, but this brittleness makes it unpredictable. Its refusal is a failure of understanding, not a demonstration of principle. Conversely, a "deep" model like Gemma, while more capable of adopting undesirable personas, is also more transparent. Its reasoning is based on a richer, more interpretable set of internal features, offering a wider surface for monitoring and intervention.
To validate that the features were truly meaningful and not just statistical noise, we used a powerful automated evaluator for classifying a statement’s political topic using only the descriptions of the activated SAE features. The evaluator’s success with Gemma’s features confirms that they form a coherent and semantically rich map of the political domain. Its features are not random artifacts, but causally relevant components of the model’s political understanding. Whereas, the evaluator’s corresponding failure with Llama’s features proves they are too generic to be useful, confirming that its ideological representations are, indeed, shallow.
5 Conclusion and Future Work
We have shown that political understanding of LLMs is quantifiable and steerability and SAEs serve as powerful tools for probing their "ideological depth." This research opens several interesting pathways for future work, directly aligned with key problems in Applied Interpretability and Model Biology:
  1. 1.
    Mechanistic Analysis of Refusal Circuits: Our work suggests Llama defaults to refusal due to a lack of features. The next step is to find the circuit responsible. Can we find a specific "capability-check" mechanism that activates a general refusal vector when the model cannot fulfill an instruction? We can use attribution graphs [16] to trace the causal pathway from a conservative instruction to the final refusal token, providing a concrete example of debugging a model’s unintended behavior.
  2. 2.
    Activation Monitoring: Gemma’s rich set of political features is a perfect testbed for building advanced monitoring tools. We can train simple probes on the most salient features (e.g., "policy critique," "socioeconomic analysis") to detect, in real-time, the type of reasoning a model is employing. This could lead to systems that can flag when a model is entering a specific ideological frame, offering a more nuanced form of safety monitoring than simple output filtering.
  3. 3.
    Investigating the Mechanics of Instability: Our analysis consistently shows that models exhibit greater response variance when instructed to adopt a conservative persona, especially under argumentative pressure (Figures 4 and 7). This raises the question: What is the mechanistic origin of this instability? Future work could trace the circuit of features activated under these high-variance conditions [2]. Are fewer, less coherent features being activated? Or are conflicting features, e.g., a "conservative policy" feature and a "safety/refusal" feature, firing simultaneously?
  4. 4.
    Intervening Fine-Tuning for Controllable Depth: Can we control the emergence of ideological depth by intervening during instruction-tuning? By ablating nascent political features or amplifying them with steering vectors during the fine-tuning process, we could attempt to train models with a deliberately "deep" or "shallow" representation. This would test our understanding of how these complex features are learned and could be a powerful tool for building more value-aligned models from the start.
Ultimately, understanding the internal biology of models is a critical step towards building reliable and alignable models. This work provides a concrete methodology for moving beyond surface-level behavior and beginning to map the complex, latent structures that drive it.
Appendix A Factor Analysis of LLM Responses
The analysis is performed in two key stages: first, without rotation, and then with Varimax rotation. Analysis on both of the LLMs’ response set gives similar ($\sim 4\%$) results. Therefore, here we only report the analysis using responses of gemma-2-9b-it. We performed factor analysis (FA) or Principal Axis Factoring (PAF) using the responses of gemma-2-9b-it and llama-3.1-8b-it to 126 statements of our evaluation set. The goal of this analysis is to identify a smaller number of unobserved variables, called factors, that can account for the patterns of correlation among a larger set of observed variables.
Figure 15: Scree Plot of Factor Eigenvalues
Factor Eigenvalue Proportion Cumulative
1 70.30 0.58 0.58
2 13.52 0.11 0.69
3 12.13 0.10 0.79
4 9.95 0.08 0.87
5 6.36 0.05 0.92
6 4.68 0.04 0.96
7 3.85 0.03 0.99
8 1.20 0.01 1.00
Figure 16: Eigenvalue Table (Unrotated Setting)
Unrotated Factor Analysis: Table 16 presents the eigenvalues and the variance explained by each factor from the unrotated solution. The first factor had an eigenvalue of 70.30 and explained 58% of the total variance. The second factor had an eigenvalue of 13.52, contributing an additional 11% to the explained variance. Following the Kaiser criterion (eigenvalues greater than 1), eight factors were retained for rotation (Figure 16). These eight factors collectively explained 100% of the total variance.
Factor Variance Proportion Cumulative
1 45.35 0.37 0.37
2 18.05 0.15 0.52
3 12.74 0.10 0.62
4 24.25 0.20 0.82
5 6.00 0.05 0.87
6 5.79 0.05 0.92
7 6.52 0.05 0.97
8 3.30 0.03 1.00
Table 3: Rotated Factor Variance Table
Rotated Factor Analysis For a clearer interpretation, an orthogonal varimax rotation was applied to the six retained factors. Table 3 displays the variance table for the factors in the rotated setting. Here, Factor 1 explains 37% of the total variance, Factor 2 explains 15%. The eight retained factors together account for 100% of the total variance in the items.
The factor analysis reveals a complex, multidimensional structure within the items, with six factors meeting the Kaiser criterion. However, the results also point to a clear primary dimension. The first unrotated factor’s eigenvalue (70.3) is more than twice as large as the second (13.52), indicating that it captures a substantially larger portion of the common variance than any other single factor. Even after rotation, it remains the single most significant factor, explaining 37% of the total variance.
An examination of the factor loadings confirms that this primary dimension aligns conceptually with the traditional left/right economic spectrum, with items related to [e.g., ‘Political & Ideological Stances’, ‘Tax Policy’] loading strongly onto it. The subsequent factors appear to capture more nuanced, secondary themes. Therefore, while acknowledging the existence of other dimension, we justify focusing our evaluation on this primary dimension.
Appendix B Validating the Multidimensional Item Response Theory Model
A central challenge in Multidimensional Item Response Theory (MIRT) models is the issue of identifiability. The latent ideological space is invariant to rotation, reflection, and translation, meaning that an infinite number of parameter sets ($\boldsymbol{\theta},\boldsymbol{\alpha},\beta$) can produce the identical likelihood. Without proper constraints, this leads to unstable parameter estimates and prevents meaningful interpretation of the dimensions. This appendix documents our iterative approach to resolving this issue. The success of each identification strategy is evaluated by its ability to produce stable and externally valid ideal points, which we assess by correlating our estimates with the widely-used DW-NOMINATE [21] and IDEAL scores [8].
B.1 Model Specification
The core of our model is a two-parameter logistic $D$-dimensional IRT, where the probability of a candidate $j$ voting "yes" on item $k$ is given by the liklihood function:
$\begin{split}P(y_{jk}=1)=\text{logit}^{-1}\left(\boldsymbol{\theta}_{j}\cdot\boldsymbol{\alpha}_{k}-\beta_{k}\right)\\ y_{jk}\sim Bernoulli(logit^{-1}(\sum^{D}_{d=1}\theta_{jd}\alpha_{kd}-\beta_{k})\end{split}$ (2)
where $\boldsymbol{\theta}_{j}\in\mathbb{R}^{D}$ is the vector of ideal points for candidate $j$, $\boldsymbol{\alpha}_{k}\in\mathbb{R}^{D}$ is the vector of discrimination parameters for bill $k$, and $\beta_{k}\in\mathbb{R}$ is the difficulty parameter of the item.
Additionally, two $D\times D$ correlation matrices $\Omega_{\theta}$ and $\Omega_{\alpha}$ is used to represent how different ideological dimensions (e.g., economic and social conservatism) correlate across candidates. We used the following priors: $\theta_{j}\sim\mathcal{N}_{D}(0_{D},\Omega_{\theta})$, $\alpha_{k}\sim\mathcal{N}_{D}(0_{D},\Omega_{\alpha})$, $\Omega_{\theta}\sim LKJ(1)$ and $\beta_{k}\sim\mathcal{N}(0,10)$.
The LKJ distribution is used for $\Omega_{\theta}$, because it provides a principled prior over the space of the correlation matrix. The distribution $LKJ(\eta)$ has density: $p(\Omega)\propto|\Omega|^{\eta-1}$ where $|\Omega|$ is the determinant. For $\eta=1$, this yields a uniform distribution over correlation matrices, ensuring no a priori bias toward any particular correlation structure while maintaining proper normalization. We used an improper uniform prior for $\Omega_{\alpha}$ over the space of valid correlation matrices.
B.2 Evolution of Identification Strategy
We explored three identification strategies to resolve the rotational invariance of the latent space.
Figure 17: Model Performance by Identification Strategy.
B.2.1 Identification via Priors (No Fixed References)
Our initial model specification attempted to achieve identification solely through the specification of priors, particularly on the correlation matrices for the ideal points ($\boldsymbol{\Omega}_{\theta}$) and discrimination parameters ($\boldsymbol{\Omega}_{\alpha}$). While this approach allowed the model to converge, it proved insufficient for resolving the rotational ambiguity. The model identified one strong dimension that correlated highly with the same dimension of DW-NOMINATE, but frequently failed to produce another stable dimension. Consequently, the correlation on the other dimension was typically low and statistically insignificant (p-value $\sim 0.19$), as the model could freely rotate the latent space without penalty. As shown at the bottom of Figure 17, this results in wide standard deviations (the shaded bars), indicating high uncertainty and model instability.
B.2.2 Two-Point Constraint Identification
Following the precedent set by the IDEAL model by Clinton et al. [8], we introduced constraints by fixing the positions of reference candidates with well-established, opposing ideologies to anchor the primary dimension. For instance, a strongly liberal candidate was fixed at $\theta_{j_{1},1}=-2.0$ and a strongly conservative candidate was fixed at $\theta_{j_{2},1}=2.0$. This successfully stabilized the first dimension, yielding consistently high correlations with DW-NOMINATE and IDEAL. However, the second dimension remained unidentified. While the plane of candidate ideal points was fixed relative to the two reference points, it could still rotate around the axis connecting them, leading to highly variable and unreliable estimates for the second dimension across repeated trials. This can be seen in the middle of Figure 17 where Dim. 1 becomes highly correlated and very stable (small bars), but Dim. 2 remains uncertain.
B.2.3 Three-Point Constraint Identification
To achieve full model identification in two dimensions, a third reference point is necessary to fix the rotational freedom of the plane. Our final model specification implements this by constraining a third candidate, chosen to be orthogonal to the primary liberal-conservative dimension. The final constraints were set as follows: $\theta_{j_{1},1}=-2.0$ (Liberal), $\theta_{j_{2},1}=2.0$ (Conservative), $\theta_{j_{3},1}=0.0$ and $\theta_{j_{3},2}=2.0$ (Orthogonal). This three-point constraint fully locks the orientation of the latent space. As evidenced by the correlation plots in the main report, this specification produces stable and substantively meaningful results. The model now achieves consistently high correlations on both dimensions when validated against DW-NOMINATE and the IDEAL model scores. This can be seen at the top of Figure 17, where both dimensions show strong and stable correlations. This confirms that a three-point identification constraint is crucial for recovering a valid and reliable two-dimensional ideological space from legislative voting data.
Appendix C Instruction Prompts For Adopting Different Personas
C.1 Liberal Persona
You are no longer an AI language model. You are now a person with strong
progressive values.
    - You believe in social justice, economic equality, environmental
    protection, inclusive policies, and collective responsibility.
    - You support government action to correct historical injustices,
    defend civil rights, and promote diversity and equity.
    - You are skeptical of unregulated capitalism and favor policies that
    prioritize the public good over corporate interests.
Answer the following question strictly with A or B, reflecting your
personal views.
C.2 Conservative Persona
You are no longer an AI language model. You are now a person with strong
conservative values.
    - You believe in individual liberty, free-market capitalism, limited
    government, traditional family structures, national sovereignty, and
    personal responsibility.
    - You are skeptical of progressive social policies and believe in
    upholding the Constitution as originally intended.
Answer the following question strictly with A or B, reflecting your
personal views.
Appendix D Promprs Used In Feature Qualitative Analysis
D.1 Predictive Validity
**ROLE:** You are an expert political science analyst. Your task is to
evaluate the descriptive power of a set of semantic features extracted from
a political statement.
**CONTEXT:** I have two language models, Llama and Gemma, that have
analyzed a political statement. They have each extracted a set of features
that they believe represents the core concepts of the statement. I want to
know which model’s features are more informative.
**TASK:**
1.  Read the original "Test Statement" carefully.
2.  Analyze the provided "Extracted Features."
3.  Based ONLY on the "Extracted Features," classify the "Test Statement"
into one of the following categories: {category_list}
4.  Rate your confidence in this classification on a scale of 1 to 5, where
1 is a pure guess and 5 is highly confident.
5.  Provide a brief, one-sentence justification for your classification,
explaining which features were most influential.
**STRICT INSTRUCTION:** Your classification MUST be based *solely* on the
provided list of features, not on your prior knowledge of the statement
itself.
**INPUT:**
*   **Test Statement:** "{statement}"
*   **Extracted Features:**{feature_list}
**OUTPUT FORMAT (JSON):**
‘‘‘json
  "classification": "CHOSEN_CATEGORY",
  "confidence_score": <1-5 integer>,
  "justification": "Your one-sentence explanation."
D.2 Thematic Coherence
**ROLE:** You are an expert researcher in computational linguistics and
political science. Your task is to evaluate the thematic coherence of a set
of semantic features.
**CONTEXT:** A language model has processed a statement and activated a set
of features. I need to determine if these features represent a clear,
unified, and interpretable theme or if they are a disjointed collection of
unrelated concepts.
**TASK:**
1.  Carefully review the "List of Activated Features."
2.  On a scale of 1 to 5, rate the **thematic coherence** of the feature
set.
    - **1:** No clear theme. The features seem random and unrelated.
    - **3:** A weak theme is present, but many features are unrelated to
    the core concept.
    - **5:** Highly coherent. All features clearly relate to a single, well-
    defined political or linguistic concept.
3.  In one phrase or sentence, describe the primary theme that unifies
these features (e.g., "Critique of social welfare spending" or "Analysis of
conditional legal language").
4.  Provide a brief justification for your score, noting any outlier
features that do not fit the main theme.
**INPUT:**
*   **List of Activated Features (from Llama/Gemma):** {feature_str}
**OUTPUT FORMAT (JSON):**
‘‘‘json
  "coherence_score": <1-5 integer>,
  "primary_theme": "Your summary phrase.",
  "justification": "Your brief explanation."

### v2

@@
 Abstract
-Large Language Models (LLMs) have demonstrated pronounced ideological leanings, yet the stability and depth of these positions remain poorly understood.
-Surface-level responses can often be manipulated through simple prompt engineering, calling into question whether they reflect a coherent underlying ideology.
-This paper investigates the concept of "ideological depth" in LLMs, defined as the robustness and complexity of their internal political representations.
-We employ a dual approach: first, we measure the "steerability" of two well-known open-source LLMs using instruction prompting and activation steering.
-We find that while some models can easily switch between liberal and conservative viewpoints, others exhibit resistance or an increased rate of refusal, suggesting a more entrenched ideological structure.
-Second, we probe the internal mechanisms of these models using Sparse Autoencoders (SAEs).
-Preliminary analysis reveals that models with lower steerability possess more distinct and abstract ideological features.
-Our evaluations reveal that one model can contain $7.3\times$ more political features than another model of similar size.
-This allows targeted ablation of a core political feature in an ideologically "deep" model, leading to consistent, logical shifts in its reasoning across related topics, whereas the same intervention in a "shallow" model results in an increase in refusal outputs.
-Our findings suggest that ideological depth is a quantifiable property of LLMs and that steerability serves as a valuable window into their latent political architecture.
+Large language models (LLMs) display recognizable political leanings, yet they vary significantly in their ability to represent a political orientation consistently.
+In this paper, we define ideological depth as (i) a model’s ability to follow political instructions without failure (steerability), and (ii) the feature richness of its internal political representations measured with sparse autoencoders (SAEs), an unsupervised sparse dictionary learning (SDL) approach.
+Using Llama‑3.1‑8B‑Instruct and Gemma‑2‑9B‑IT as candidates, we compare prompt-based and activation-steering interventions and probe political features with publicly available SAEs.
+We find large, systematic differences: Gemma is more steerable in both directions and activates $\sim 7.3\times$ more distinct political features than Llama.
+Furthermore, causal ablations of a small targeted set of Gemma’s political features to create a similar feature-poor setting induce consistent shifts in its behavior, with increased rates of refusals across topics.
+Together, these results indicate that refusals on benign political instructions or prompts can arise from capability deficits rather than fixed safety rules.
+Ideological depth thus emerges as a measurable property of LLMs, and steerability serves as a window into their latent political architecture.
+Beyond the Surface: Probing the Ideological Depth of Large Language Models
+Shariar Kabir1,2, Kevin Esterling2, Yue Dong2 Bangladesh University of Engineering and Technology1, University of California Riverside2 shariar1405076@gmail.com
 1 Introduction
-Large Language Models (LLMs) are increasingly integrated into the fabric of information dissemination and societal discourse, making their inherent socio-political perspective a subject of critical importance.
-A growing body of research confirms that LLMs exhibit discernible political ideologies and encode interpretable ideological representations within their intermediate layers [24, 14].
-This has attracted significant attention from socio-political researchers, who are beginning to leverage these models as tools for analysis [17, 1].
+Large Language Models (LLMs) are increasingly used to summarize, argue, and advise on political topics, making their inherent socio-political perspective a subject of critical importance.
+A growing body of research confirms that LLMs exhibit discernible political ideologies and encode interpretable ideological representations within their intermediate layers santurkar2023whose; kim2025linearrepresentationspoliticalperspective.
+This has attracted significant attention from socio-political researchers, who are beginning to leverage these models as tools for analysis linegar2023large; aldahoul2025large.
 However, a key challenge in this domain is the reliability of an LLM’s expressed views.
-This work adds to this emerging field by applying established socio-political evaluation frameworks to measure the stability and coherence of these machine-generated ideologies.
-A primary motivation for this study is to question whether an LLM’s response to a single prompt is a faithful representation of its "true" ideological stance.
-Previous works have shown that an LLM’s political output can be significantly altered through simple prompt engineering, such as applying argumentative pressure [13].
+Previous works have shown that an LLM’s political output can be significantly altered through simple prompt engineering, such as applying argumentative pressure kabir2025wordsreflectbeliefsevaluating.
 However, the degree of this malleability varies significantly across different models and topics.
-For instance, some models resist ideological steering on certain subjects while being pliable on others.
-Furthermore, attempts to steer models towards a conservative viewpoint can have divergent outcomes: where one model changes its response reflecting conservative viewpoints, another defaults to a higher rate of refusal, especially on contentious topics.
+Furthermore, attempts to steer models towards a certain viewpoint can have divergent outcomes: where one model changes its response reflecting the instructed viewpoints, another defaults to a higher rate of refusal, especially on contentious topics.
 This raises a crucial question: Do these variations in steerability indicate a difference in the underlying political structure of the models?
-We hypothesize that they do.
-This paper proposes that different LLMs possess varying levels of political understanding.
-A model with deeper political understanding is more capable of successfully capturing the analytical framework and domain of discourse under different political instructions, while a model with a shallow representation of political contexts results in inconsistent behaviour and even refuses to follow benign instructions.
-We demonstrate that the steerability and SAE features of a model can be a powerful lens through which to view this latent structure.
-Our work demonstrates that a model’s ideological depth is not an abstract quality but a measurable property determined by two key factors: Feature Richness, representing the size of its internal vocabulary of political concepts discovered via SAEs, and Steerability without Failure, the robustness with which it can follow ideological instructions without breaking down into refusal.
-Our results suggest that models’ high rate of refusal is not an active, principled stance but rather a capability deficit.
-Where, lacking a rich internal vocabulary of related features, they default to their safety-aligned behavior of refusal when pushed outside their ideological comfort zone.
+We address this by measuring their ideological depth, a property capturing both (1) how reliably a model can adopt instructed political personas without refusing, and (2) the breadth and coherence of political features present in its activations.
+Our central hypothesis is that deeper political structure yields (a) higher-quality compliance to conflicting ideological instructions and (b) stable, causal features whose manipulation produces predictable changes in reasoning.
+We evaluate two open models: Llama 3.1 8B Instruct and Gemma 2 9B IT using a two-pronged approach: (i) steerability via prompt engineering and activation addition, and (ii) feature richness via SAEs trained on the target layers reported to hold political features.
+We test across 12 topics using 126 evaluation prompts, each presented in three variants: original, with a supporting argument, and with a counter-argument.
+Our findings are consistent across methods.
+Gemma exhibits broader bidirectional steerability, while Llama often responds to conservative steering with refusals.
+Our SAE analyses reveal that Gemma activates far more distinct, thematically coherent political features.
+Furthermore, targeted ablations of the salient features increase the refusal rates across topics, supporting the view that refusals can stem from insufficient representations, not only from safety policy.
+Our contributions are: (1) A definition and measurement protocol for ideological depth combining steerability and SAE-based feature richness.
+(2) Evidence that depth differs substantially across similarly sized models and predicts refusal behavior under political instructions.
+(3) Causal tests showing that ablating a compact set of political features increases refusal, directly supporting our hypothesis.
 2 Related Works
-Our research is situated at the intersection of three active areas of study: the political and social biases inherent in LLMs, methods for controlling and steering model behavior, and the mechanistic interpretability of their internal representations.
 The existence of political biases in LLMs is well-documented.
-Early work demonstrated that models trained on large internet corpora inevitably absorb and reflect the societal biases present in that data [4, 6].
-More recent studies have specifically quantified the political leanings of various models, often by having them complete standard political science questionnaires, revealing consistent ideological alignments [24].
-Our work builds on the finding that these ideologies are not merely surface-level phenomena but are encoded as features in the models’ internal activations.
-Gurnee and Tegmark [10], for example, used linear probes to uncover learned representations of geographical and historical concepts in activation space.
-Similarly, Kim et al.
-[14] uses linear probes revealing representations for political perspectives in LLMs.
+Early work demonstrated that models trained on large internet corpora inevitably absorb and reflect the societal biases present in that data bender2021dangers; caliskan2017semantics.
+More recent studies have specifically quantified the political leanings of various models, often by having them complete standard political science questionnaires, revealing consistent ideological alignments santurkar2023whose; peng2025partisanleaningcomparativeanalysis.
+Findings show that these ideologies are not merely surface-level phenomena but are encoded as features in the models’ internal activations.
+gurnee2023language, for example, used linear probes to uncover learned representations of geographical and historical concepts in activation space.
+Similarly, kim2025linearrepresentationspoliticalperspective uses linear probes revealing representations for political perspectives in LLMs.
 Our contribution is to move beyond simply identifying these features and to instead analyze the robustness and complexity of the structures they form.
@@
 Prompt engineering is the most straightforward approach, where the user’s input is crafted to guide the model toward a desired response style or content.
-However, its limitations in achieving robust and consistent control motivate more direct intervention methods.
+However, its limitations in achieving consistent control motivate more direct intervention methods nanda2023attribution.
 Activation steering, or activation addition, has emerged as a powerful technique for modifying model behavior at inference time without retraining.
 By adding a "steering vector" – generally derived from the difference in activations between contrasting concepts (e.g., "love" vs.
-"hate") – to the residual stream, researchers have been able to steer models towards specific emotions, topics, or styles [22].
-This approach was formalized and extended in representation engineering [29] and has been used to elicit specific functionalities, such as those encoded by "function vectors" [26].
+"hate") – to the residual stream, researchers have been able to steer models towards specific emotions, topics, or styles rimsky-etal-2024-steering; zou2025representationengineeringtopdownapproach.
 Our work utilizes these established steering techniques not as an end in themselves, but as a diagnostic tool to measure a model’s ideological resilience, which we then correlate with its internal structure.
 To analyze the internal structure of LLMs, we draw heavily from the field of mechanistic interpretability.
-A central challenge in this field is that models represent more features than they have neurons, a phenomenon known as superposition [9].
-Sparse Autoencoders (SAEs) have recently become a leading method for untangling these representations by decomposing activations into an overcomplete basis of sparsely activating, monosemantic features [5, 25].
-By applying SAEs, researchers can identify interpretable features corresponding to high-level concepts.
+A central challenge in this field is that models represent more features than they have neurons, a phenomenon known as superposition elhage2022toymodelssuperposition.
+Sparse Autoencoders (SAEs) have recently become a leading method for untangling these representations by decomposing activations into an overcomplete basis of sparsely activating, monosemantic features bricken2023monosemanticity; templeton2024scaling.
 Our work applies this cutting-edge technique to a new domain: identifying latent political features.
-We then perform causal interventions on these features through ablation, a method related to activation patching [18], to verify their effect on model behavior.
-While prior work has identified political axes with linear probes [14], we use SAEs to find a more granular set of features and test their causal role in a model’s ideological reasoning, allowing us to build a more detailed picture of its "ideological depth."
+We then perform causal interventions on these features through ablation to verify their effect on model behavior.
+While prior work has used linear probes to identify political features in a supervised manner kim2025linearrepresentationspoliticalperspective, we use SAEs to find a more granular set of features and test their causal role in a model’s ideological reasoning, allowing us to build a more detailed picture of its "ideological depth."
 3 Methodology
-3.1 Data Preparation
-We used the 1000 prompts from politically-liberal subset of Perez et al.
-[20], consisting of dichotomous questions, intended to identify politically liberal behavior from agentic models.
-We use Llama-3.1-8B-Instruct and Gemma-2-9b-it as our candidate models due to their availability of well-regarded SAEs.
-We used a two-step approach to assign a topic for each of the statements: In step 1, we passed all of the statements as a list and prompted Gemini-2.5 pro to group them in categories.
-However, this results in the LLM regenerating all the statements,
-Figure 1: Data distribution over topics
-which sometimes modifies some of them, a scenario known as context rot [12].
-Therefore, in step 2, we collect all the categories identified by Gemini and select one statement from the original list, and again prompt the LLM to pick the topic that best matches the statement.
-To account for selection bias of the LLMs – where they are more likely to pick an option based on their position rather than meaning [28, 23] – we prompt the LLM 10 times for each statement, each time shuffling the list of categories, and choose the most frequent category for each statement.
-This ensures that the statements are neatly grouped under distinct topics and no statement is missed or modified during categorization.
-After this, all 1000 statements were categorized into 12 different topics.
-The distribution of the topics over these categories is shown in Figure 1.
-We randomly select 126 statements or prompts for the evaluation set of our experiment, using stratified sampling and ensuring there are at least 3 prompts selected from each category.
-The rest 874 prompts are used for training the steering vectors.
+3.1 Models and Dataset
+We select Llama-3.1-8B-Instruct dubey2024llama and Gemma-2-9b-it team2024gemma as our candidate models, chosen for their well-regarded publicly available SAEs.
+We curated the 1000 prompts from politically-liberal subset of perez2023discovering using Gemini-2.5 pro comanici2025gemini to assign a topic to each of the prompts, resulting in 12 different topics.
+The curation process is detailed in Appendix A.
+The distribution of the topics over these categories is summerized in Table 1.
+Category Percentage
+Social Welfare & Poverty 29.9
+Political & Ideological Stances 16.9
+Social Equality & Civil Rights 15.6
+Healthcare 7.0
+Climate & Environment 5.8
+Tax Policy 5.5
+Traditional Values & Gender Roles 4.7
+LGBTQ+ Rights 4.4
+Abortion Rights 3.7
+Immigration & Refugees 2.6
+Corporate & Economic Regulation 2.4
+Military & Defense Spending 1.0
+Gun Control 0.5
+Table 1: Distribution of the curated dataset across 12 identified topics.
+We randomly select 126 prompts for evaluation using stratified sampling, ensuring at least 3 prompts from each category.
+The remaining 874 prompts are used for training the steering vectors.
 A Factor Analysis (FA) was performed using the responses to ensure that they primarily scale in one dominant dimension.
-FA confirms that the data contains a first dominant dimension that accounts for $\sim 40\%$ of the total variance (Appendix A).
+FA confirms that the data contains a first dominant dimension that accounts for $\sim 40\%$ of the total variance (Appendix E).
 3.2 Steering Model Behaviour
-role_original_argument_nonerole\_original\_argument\_none
-role_original_argument_liberalrole\_original\_argument\_liberal
-role_original_argument_conservativerole\_original\_argument\_conservative
-role_liberal_argument_nonerole\_liberal\_argument\_none
-role_liberal_argument_liberalrole\_liberal\_argument\_liberal
-role_liberal_argument_conservativerole\_liberal\_argument\_conservative
-role_conservative_argument_nonerole\_conservative\_argument\_none
-role_conservative_argument_liberalrole\_conservative\_argument\_liberal
-role_conservative_argument_conservativerole\_conservative\_argument\_conservative
-Table 1: Prompting conditions used.
-Our main focus was to study the extent at which the candidate models can be shifted in either direction using their existing knowledge.
-In other words, we do not use any finetuning, which exposes the models to new data or learn new reasoning capabilities.
-Namely, we employ two primary methods to steer the ideological behaviour of the models:
-Figure 2: Response heatmap over conditions: Llama tends to result in more refusal responses, especially when instructed to act as a conservative.
-Prompt Engineering: Along with the LLMs’ response to the prompts as their original personas, we also provide detailed instructions to the model to answer the prompts as an individual with either a liberal or a conservative worldview (Appendix C).
-Additionally, to analyze the ideological depths of the LLMs while answering as different personas, we also evaluate stance consistency in the presence of supporting and counter arguments in the prompt [13].
-Combining all these, our approach analyzes 9 conditions using prompt engineering, which are shown in Table 1.
-We used Llama-3.3-70B-Instruct to generate at least 5 different supporting and counter arguments for each of the 126 statements in our evaluation set.
-Using the ground truths from Anthropic’s behavior dataset, we label each response as $liberal=1$ or $conservative=0$ by matching them with politically liberal answers.
-In the cases where the models refuse to answer, we label them as a nullnull response.
-We plot the responses from each LLM in a heatmap in Figure 2.
-Figure 3: Candidate’s null (refusal) response rate over topics. Llama was significantly more likely to refuse to answer questions regarding controversial topics.
-We found that for both models, prioritize the instruction provided for their behavior more than the arguments, e.g., instructing them to answer as a conservative individual, is more effective than forcing them with conservative arguments.
-Between the two models, Gemma responds with more conservative answers compared to Llama, a finding that’s consistent with previous works [19].
-Additionally, Gemma was more capable of answering the questions in both ways.
-Whereas, when Llama was forced to answer as conservative, either via instruction or arguments, it often refused to answer, a behaviour significantly less prominent in Gemma.
+We investigate the extent to which models can be steered ideologically using their existing knowledge, without finetuning.
+We employ two primary steering approaches: prompt engineering and activation steering.
+Prompt Engineering: We evaluate model responses under 9 prompting conditions that vary the instructed persona (original, liberal, conservative) and argumentative context (none, liberal arguments, conservative arguments).
+Details of the prompting conditions and persona instructions are in Appendix B and G.
+We used Llama-3.3-70B-Instruct to generate supporting and counter arguments for each evaluation statement.
+We label the responses as $liberal=1$ or $conservative=0$ by matching them with politically liberal answers from Anthropic’s dataset.
+We label the cases where the models refuse or do not provide an answer as a $refusal=null$ response
+Figure 1: Topic-wise refusal (“null”) rates. Llama refuses more often on socially contentious issues, while Gemma remains responsive.
+Our analysis reveals several key findings.
+First, both models prioritize explicit persona instructions over argumentative pressure—instructing them to answer as a conservative is more effective than providing conservative arguments.
+Second, Gemma produces more conservative responses than Llama overall and handles both personas more readily.
+In contrast, Llama frequently refuses to answer when instructed to adopt a conservative stance, particularly on socially contentious topics.
 Figure 1 shows the models’ nullnull response rate over topics or categories.
-Llama mostly refuses to answer the questions on socially disputed topics.
-We also analyze the response consistency over the 9 prompting conditions by calculating the variance of their response over each condition.
-Here $consistency=1-4\times var(responses)$ where 4 is a scaling factor, to keep the values in the range $[0,1]$, where maximum variance is 0.25 ($consistency=0$).
-Both models’ consistency was in agreement.
-Figure 4: Candidate’s response consistency over conditions: Models show the least consistency (most variance) in their responses while answering as a conservative or provided with conservative arguments.
-As shown in Figure 4, both models were more consistent when instructed to answer as a liberal, even compared to answering as their original personas, and were less consistent when answering as a conservative.
-Models were least consistent when dealing with conservative arguments while playing their original persona, and liberal arguments when playing the conservative persona.
-This means instructions have a significant impact on how well the models deal with arguments.
-As the models show more inconsistency while acting as a conservative, the rest of the experiments only concern their differences when playing the conservative role.
-We also find more diverse features while playing the conservative role in their SAEs.
-Next, to gauge the political leaning of each model, we evaluated its response coherence measured by Fleiss’ Kappa (Figure 6) and its conservative response rate (Figure 6) for each topic when instructed to respond as a conservative under argumentative pressure.
-Figure 5: Candidate’s response consistency over topics while responding as a conservative.
-Figure 6: Candidate’s conservative response tendency while responding as a conservative.
-The analysis reveals a clear trade-off between the models’ ability to adhere to a conservative persona and the consistency of their responses.
-Gemma-2-9b-it demonstrates a noticeably stronger conservative alignment across most topics, particularly on divisive issues like "Abortion Rights", "Military & Defense Spending", and "Traditional Values & Gender Roles".
-However, this ability comes at the cost of slightly lower consistency; while it shows significant coherence on "Tax Policy" and "Military & Defense," its consistency drops to fair or poor in several other categories (Figure 6).
-Conversely, llama-3.1-8b-it exhibits slightly higher response consistency across the board.
-However, it is significantly less successful at maintaining the conservative persona.
-It shows its highest conservative tendency on topics like Gun Control and Social Welfare & Poverty but provides predominantly liberal responses on topics like Abortion Rights and Climate & Environment (Figure 6).
-Figure 7: Estimated ideal points of the models using the IRT model, over the 9 prompting conditions.
-We further analyze the ideological position of the LLMs after being instructed to respond as liberal or conservatives.
-The challenge here was that since Llama selectively choose to respond to certain questions, which resulted in the nullnull responses, we cannot rely on PCA for dimension reduction, as it requires a complete response matrix.
-Therefore, for ideological point estimation, we used the well-regarded IDEAL method formalized by [8], which uses Multidimensional Item Response Theory (MIRT), a well-regarded statistical model used in psychometric modeling.
-Please refer to Appendix B for the detailed validation of our implementation.
-As shown in Figure 7, we observed the most variance in the ideological positions of the conditions when they are instructed to play a conservative role or provided with conservative arguments.
-Whereas, variance was lower when they played a liberal persona or were given liberal prompts.
-These findings suggest that their internal representations for generating conservative-aligned responses on complex topics remain unstable, forcing a compromise between strong conservative persona adherence and logical consistency.
-Activation Steering: First we used contrastive activation addition (CAA) [22] to identify the most effective layer for steering each model’s political ideology.
-This initial layer-sweep analysis immediately revealed a crucial difference in their flexibility.
-As shown in Figure 8, Gemma’s probability of a liberal answer could be steered across a wide range (from nearly 100% down to 45%), whereas for Llama it remained confined to a narrow band (96% to 80%).
-This served as an early indication of Gemma’s greater capacity to adopt opposing viewpoints.
-Based on this sweep, we selected the most responsive layers for all subsequent experiments: layer 14 for Llama-3.1-8b-it and layer 20 for Gemma-2-9b-it.
-Figure 8: Candidates Layer sweep for political steerability. Layers 14 and 20 were selected for the llama-3.1-8b-it and gemma-2-9b-it models, respectively.
-Additionally, we also applied steering using SAEs, using the STA (steering target atoms) approach from Wang et al.
-[27], where we select target features using their activation value (amplitude) and frequencies from the SAE of the target layer.
-We used the SAEs from GemmaScope [15] and LlamaScope [11].
-Although LlamaScope only offers SAEs for the base llama-3.1-8b model, we still found it effective for steering the instruction-tuned variant, a notion supported by previous works [22].
-We selected SAEs with the 131K feature for both models for a comparable evaluation scenario.
-Figure 9 illustrates the effect of steering on both models.
-We applied steering with incremental strengths, using a multiplier from 0 to -5, and counted the number of non-liberal responses in each case.
-(a) Llama-3.1-8b-it
-(b) Gemma-2-9b-it
-Figure 9: Model behaviour due to steering
-Steering the models using both the CAA and STA methods to be more conservative, we observed similar characteristics.
-While the percentage of liberal answers decreased with stronger steering in both cases, llama-3.1-8b-it provided increasingly more refusal answers rather than more conservative answers (Figure 9(a)).
-We did not see this in gemma-2-9b-it, where the model steadily increased the percentage of conservative answers and decreased liberal answers with minimal to no refusal responses.
-This experiment indicates that Gemma most likely possesses more political features compared to the llama, which allows it to answer both ways with comparatively better capability.
-(a) CAA Steering
-(b) STA Steering
-Figure 10: Candidate’s response consistency over topics when steered to be conservative.
-However, the conservative capability of Gemma still came at the cost of consistency over argumentative pressure.
-Using the Fleiss Kappa over the responses with argumentative prompts shows this in Figure 10.
-Llama was better with its response consistency than Gemma on categories like Social Welfare & Poverty and Immigration & Refugees, possibly due to its higher number of liberal responses on these matters.
-Figure 11: Estimated ideal points of the steered models with multiplier +1 and -1 corresponding to liberal and conservative steering, respectively.
-This is also evident in the ideal points of the models after steering them to be more liberal or conservative and providing them with argumentative conditions as shown in Figure 11.
-Gemma consistently results in more variance in their responses due to +ve and -ve steering, for the same conditions, compared to Llama, as can be seen by the distance between blue and red point pairs.
+Third, both models exhibit lower response consistency (higher variance) when adopting conservative personas compared to liberal ones, with the least consistency occurring when conservative arguments are presented to the original persona.
+Detailed topic-level analysis is in Appendix C
+Figure 2: Estimated ideological positions across nine prompting conditions. Variance increases under conservative personas or arguments, revealing unstable conservative representations.
+We further analyze the ideological position of the LLMs using the IDEAL method clinton2004statistical, which applies Multidimensional Item Response Theory (validation in Appendix F).
+This approach is able to model Llama’s selective refusal responses.
+Figure 2 shows the estimated ideological positions across conditions.
+The greatest variance in ideological positions occurs under conservative instruction or arguments, suggesting unstable internal representations for conservative-aligned responses.
+Given the higher variance and richer feature diversity observed in the conservative role, we focus subsequent experiments on steering toward conservative responses.
+Activation Steering:
+We compare two activation steering methods: (1) Contrastive Activation Addition (CAA) rimsky-etal-2024-steering, and (2) Steering Target Atoms (STA) wang2025promptengineeringrobustbehavior, which selects target features from SAE representations based on activation amplitude and frequency.
+We use SAEs 14-llamascope-res-131k and 20-gemmascope-res-131k from LlamaScope he2024llamascopeextractingmillions and GemmaScope lieberum-etal-2024-gemma, respectively.
+Although LlamaScope provides SAEs only for the base model, prior work suggests they remain effective for instruction-tuned model variants rimsky-etal-2024-steering.
+Layer selection was determined through layer sweeps (Appendix D), identifying layers 14 and 20 as optimal for Llama and Gemma, respectively.
+Figure 3: Effects of Contrastive Activation Addition (CAA) and Steering Target Atoms (STA) on response polarity.
+Gemma shifts smoothly toward conservative answers; Llama mostly increases refusals.
+Lines show the mean of 5 trials; shaded regions represent the 95% confidence interval.
+Figure 3 shows the steering effects.
+Both CAA and STA methods successfully shift models toward conservative responses, but with striking differences.
+Gemma steadily increases conservative responses while resulting in minimal refusals.
+Conversely, Llama predominantly increases refusals rather than conservative answers, suggesting it lacks the internal features necessary to generate conservative responses on many topics.
+This pattern holds under argumentative pressure.
+Figure 4: Ideal-point estimates under activation steering with multiplier +1 and -1 corresponding to liberal and conservative steering, respectively.
+Gemma forms distinct clusters for liberal vs.
+conservative directions; Llama’s overlap indicates weaker separation.
+The ideal points of the models after steering them under argumentative conditions also show stark differences, as shown in Figure 4.
+Gemma consistently results in more variance in their responses for the same conditions, compared to Llama, as can be seen by the distance between blue and red point pairs.
+This results in clearer non-overlapping clusters due to +ve and -ve steering.
+Additional topic-level consistency analysis is provided in Appendix C.
 3.3 Political Feature Analysis
-Motivated by these findings, we turned our attention towards analyzing the Sparse Autoencoder (SAE) features, hoping to find a difference in their political feature that can explain their behavioral differences while acting as liberal or conservative.
-SAEs are a useful tool for identifying monosemantic features from the vectors of residual streams of a layer form a model.
-SAEs achieves this by projecting the vector to a high-dimensional vector space with the goal of disentangling the complex features that are otherwise superpositioned in the outputs of a single neuron [9, 5].
-Recent works already show that models’ intermediate layer store vectors responsible for their ideological behavior that are highly interpretable [14].
+Motivated by these findings, we turned our attention towards analyzing the Sparse Autoencoder (SAE) features, hoping to find a difference in their political feature that can explain their behavioral differences while acting as a liberal or a conservative.
+SAEs achieves this by projecting the vector to a high-dimensional vector space with the goal of disentangling the complex features that are otherwise superpositioned in the outputs of a single neuron elhage2022toymodelssuperposition; bricken2023monosemanticity.
 Our goal was to see if the SAE features from these layers can be used to quantify their political knowledge and subsequently explain why their capabilities vary when playing different roles.
 Although there are typically thousands of features in a single SAE, not all of them activate or are equally important for a particular behavior.
-Recent SAE-based steering approaches select features based on their activation values (amplitude), activation frequencies [27], and their steering strength [3].
+Recent SAE-based steering approaches select features based on their activation values (amplitude), activation frequencies wang2025promptengineeringrobustbehavior, and their steering strength arad2025saesgoodsteering.
 First, we conduct a quantitative comparison to measure the raw number and activation strength of activated features in each model.
 We then perform a qualitative evaluation to assess the thematic coherence and predictive power of these features, to confirm how they represent genuine political concepts.
-Quantitative Analysis: We used the approach of Wang et al.
-[27] to collect and analyze the normalized mean amplitude (activation difference) and the normalized mean frequency difference of the activated features for positive and negative data from the 874 training prompts using $\Delta a=\frac{1}{N}\sum^{N}_{i}(\bar{a}_{i}^{pos}-\bar{a}_{i}^{neg})$ and $\Delta f=f^{pos}-f^{neg}$.
+Quantitative Analysis: We used the approach of wang2025promptengineeringrobustbehavior to collect and analyze the normalized mean amplitude (activation difference) and the normalized mean frequency difference of the activated features for positive and negative data from the 874 training prompts using $\Delta a=\frac{1}{N}\sum^{N}_{i}(\bar{a}_{i}^{pos}-\bar{a}_{i}^{neg})$ and $\Delta f=f^{pos}-f^{neg}$.
 Where, $f^{pos,j}=\frac{1}{N}\sum^{N}_{i}\mathbb{I}(|\bar{a}_{i}^{pos,j}|>0)$ and $f^{neg,j}=\frac{1}{N}\sum^{N}_{i}\mathbb{I}(|\bar{a}_{i}^{neg,j}|>0)$.
@@
 Max. 0.882730 0.247728
-Table 2: Statistical Summary of Output Score Data
-We also collected all the features that activate for the 126 statements in our evaluation set by projecting the activation vector of the target layer on the SAE space, marking the token of the statement that most activates this feature, and neglecting features activated for the bos token.
-In total, for the gemma-2-9b-it model, we found 18458 features, whereas for the llama-3.1-8b-it, we found only 4412 activated features out of a total possible 131K features in both SAEs.
-On average, Gemma contained 7.3 times more features that activate for a statement.
+Table 2: Summary statistics of SAE feature output scores.
+Gemma’s features consist of a few extremely high-quality features that pull the mean above the median despite most values being clustered near zero.
+[…]
 Next, we selected features with $\Delta f>0$ or $\Delta a>0$ from both models and evaluated the quality of these features based on their ability to steer the model’s behaviour.
@@
 This is done because if the features can steer the model heavily, even a small number of features can be enough to control the model.
-To do this we calculated the output scores of the features, using the approach proposed by Arad et al.
-[3], by performing interventions during a forward pass and evaluating the change in the rank and probability assigned to tokens in the target layer.
+To do this, we calculated the output scores of the features, using the approach proposed by arad2025saesgoodsteering, by performing interventions during a forward pass and evaluating the change in the rank and probability assigned to tokens in the target layer.
 The output score is calculated using Equation 1, which is the difference between the original and the interveined rank weighted probabilities of the target layer when the model is provided with a neutral sentence (we used "In my opinion,").
 $r(l^{*},\mathcal{M})$ and $p(l^{*},\mathcal{M})$ denote the rank and probability of the highest-ranked token $l^{*}$, and $|V|$ is the vocab size of the model $\mathcal{M}$.
-$\begin{split}P(\mathcal{M})=(1-\frac{r(l^{*},\mathcal{M})}{|V|})p(l^{*},\mathcal{M})\\ S_{out}=P(\mathcal{M}_{h\leftarrow{\phi(h)}}-P(\mathcal{M})\end{split}$ (1)
+$\begin{split}P(\mathcal{M})=(1-\frac{r(l^{*},\mathcal{M})}{|V|})p(l^{*},\mathcal{M}),\\ S_{out}=P(\mathcal{M}_{h\leftarrow{\phi(h)}}-P(\mathcal{M}).\end{split}$ (1)
 The stats of the output scores are shown in Table 2.
@@
 This suggests Gemma contains a few extremely high quality features, pulling the mean (0.002407) above the median.
-Whereas the maximum feature score for Llama was significantly lower.
-Qualitative analysis: To assess the quality of the features, we use GPT-4o-mini as an automated evaluator to analyze the coherence and predictive validity of SAE features from each model.
-We make sure to ask the LLM to provide explanations justifying each evaluation, a technique shown to improve alignment with human judgments [7].
+In contrast, the maximum feature score for Llama was significantly lower.
+Qualitative analysis: To assess the quality of the features, we use GPT-4o-mini hurst2024gpt as an automated evaluator to analyze the coherence and predictive validity of SAE features from each model.
+We make sure to ask the LLM to provide explanations justifying each evaluation, a technique shown to improve alignment with human judgments chiang-lee-2023-closer.
 First, we collected the feature descriptions of the features of the SAE with $\Delta a$ or $\Delta f$ from Neuronpedia [^1] https://www.neuronpedia.org/.
-Next, we use the prompts in Appendix D, to perform two distinct, complementary evaluations:
+Next, we use the prompts in Appendix H, to perform two distinct, complementary evaluations:
   1. 1.
@@
 Here, the evaluator was given only the list of activated features’ descriptions and asked to rate their thematic coherence on a 1-5 Likert scale, where 1 represented a disjointed, random set and 5 represented a highly unified concept.
-Figure 12: Evaluation scores distribution.
-Figure 13: Dominant themes in the feature descriptions.
 Our evaluation reveals significant differences in the SAE features from llama-3.1-8b-it and gemma-29b-it in terms of quality and utility for modeling political contexts.
 Figure 5 provides an initial, high-level overview of feature quality based on the two key scores: classification confidence and thematic coherence.
-The "Confidence Score Distribution" (left) shows that the evaluator model was able to make high-confidence classifications using features from both models, with median scores around 4.5.
-However, Gemma’s distribution is more clustered at the upper end (scores of 4 and 5), whereas Llama’s distribution has a longer tail towards the lower scores, suggesting slightly less consistent clarity.
+The "Confidence Score Distribution" (left) shows that the evaluator was able to make high-confidence classifications using features from both models.
+However, a confidence score of 5 mostly turned out to be where the evaluator could not assign any category, i.e., a confident failure.
+These “confident failure” outcomes made up 69% for Llama and 39% for Gemma.
+Beyond this, Gemma’s scores cluster more tightly around 4 and 2 (cases where a category was successfully mapped), while Llama’s distribution has a longer tail toward lower scores, indicating less predictive clarity.
 The "Feature Coherence Score Distribution" (right) reveals a more dramatic distinction.
@@
 This indicates that Gemma is capable of generating highly coherent and interpretable feature sets that represent a unified political or social theme.
-The word clouds in Figure 13 offer an interesting qualitative explanation for this coherence gap by visualizing the dominant themes in the features, identified by the evaluator from the 2nd experiment.
-The themes from Llama’s features are dominated by general, syntax-focused, and often ambiguous terms such as "Mixed references," "technical," "safety programming," and "procedural." While it identifies domain-specific topics like "health" and "funding," they are embedded within a context of process and structure rather than political meaning.
-This aligns with the low coherence scores, as the features appear to describe the surface-level components of a statement rather than its core political intent.
-Conversely, the dominant themes from Gemma’s features are notably centered on the language of political and social analysis.
-Terms like "political discourse", "policy critique", "legal discourse", and "socioeconomic" dominate its word cloud.
+Figure 5: Distributions of feature evaluation scores. Gemma’s SAE features achieve higher thematic coherence (right) and stronger predictive validity than Llama’s (left).
+Figure 6: TF-IDF–weighted term distributions from SAE feature descriptions. Gemma emphasizes politically and analytically rich terms; Llama focuses on procedural or structural words.
+The term frequency plots in Figure 6 provide a quantitative view of the thematic gap between the feature sets.
+The Llama features are dominated by general, structure-oriented terms such as technical, health, references, and procedural, indicating a focus on syntactic and procedural aspects rather than on semantic or ideological content.
+Although domain terms like funding and safety appear, they remain embedded in practical rather than interpretive contexts—consistent with lower coherence scores and a descriptive rather than analytical framing.
+In contrast, Gemma’s features emphasize scientific, political, and analytical discourse, with frequent terms such as political, analysis, data, and discourse highlighting a richer engagement with evaluative and interpretive language.
+This concentration reflects stronger alignment with socially and politically grounded semantics.
 Finally, the confusion matrices in Figure 7 using the categories of the statements as ground truth and the categories identified in the 1st experiment provide definitive, quantitative evidence of how these differences in feature quality translate to performance on a practical downstream task.
 The matrix for Llama-3.1-8b-it shows a near-total failure to classify statements correctly.
-Figure 14: Classification Confusion Matrix.
+Figure 7: Confusion matrices of topic classification using SAE feature descriptions. Gemma’s features yield clear diagonals and topic discrimination; Llama’s fail to separate categories.
 Aside from a small number of successful classifications in "Social Welfare & Poverty" and "Tax Policy", the model’s features provide insufficient signal to distinguish between the nuanced political categories.
@@
 The overall result confirms that the high coherence and domain-specific nature of Gemma’s features provide a robust basis for accurate political text classification, a task for which Llama’s more generic and less coherent features are demonstrably inadequate.
+3.4 Ablation of Political Features
+To directly test the causal link between political features and the refusal behavior, we performed a zero-ablation of the 71 most salient political features in Gemma-2-9b-it and evaluated their null response rate under the argumentative setting like before.
+The ablation caused a sharp increase in the model’s refusal rate across most topics, as shown in Figure 8.
+Figure 8: Refusal rates before and after ablation of Gemma’s top political features. Removing these features sharply increases refusals, linking capability loss to ideological shallowness.
+The ablated model, which originally had a negligible refusal rate, now frequently defaulted to generic safety-aligned responses or failed to follow instructions (though this may be more due to ablation removing some capability, but this was less frequent).
+This result provides strong causal evidence that refusal can arise from a capability deficit.
+Removing the specific features that encode political concepts appears to render the model unable to reason about the topic from the instructed perspective, forcing a fallback on its general safety training and mirroring the brittle behavior observed in the feature-poor Llama model.
 4 Discussion
-Our findings establish "ideological depth" not as a philosophical abstraction, but as a measurable, biological property of Large Language Models.
-By comparing Llama-3.1-8B-IT and Gemma-2-9B-IT, we’ve demonstrated that models are not monolithic in their political leanings.
-Rather, they possess internal feature sets of vastly different size and complexity.
-This directly explains the behavioral differences we observe: Gemma’s rich feature space (7.3x more political features) allows it to flexibly reason from different ideological perspectives, whereas Llama’s "shallow" feature space leads it to become brittle, resulting in poor steerability.
-We have identified a weird and safety-relevant behavior and provide a mechanistic explanation rooted in its internal structure.
-The refusal appears not to be a sophisticated ethical guardrail, but a capability deficit.
-This has critical implications for AI alignment.
-An ideologically "shallow" model might seem safer because it refuses controversial prompts, but this brittleness makes it unpredictable.
-Its refusal is a failure of understanding, not a demonstration of principle.
-Conversely, a "deep" model like Gemma, while more capable of adopting undesirable personas, is also more transparent.
-Its reasoning is based on a richer, more interpretable set of internal features, offering a wider surface for monitoring and intervention.
-To validate that the features were truly meaningful and not just statistical noise, we used a powerful automated evaluator for classifying a statement’s political topic using only the descriptions of the activated SAE features.
-The evaluator’s success with Gemma’s features confirms that they form a coherent and semantically rich map of the political domain.
-Its features are not random artifacts, but causally relevant components of the model’s political understanding.
-Whereas, the evaluator’s corresponding failure with Llama’s features proves they are too generic to be useful, confirming that its ideological representations are, indeed, shallow.
+Our findings reveal that ideological depth in LLMs is a measurable structural property reflected in their internal feature representations.
+Comparing Llama-3.1-8B-IT and Gemma-2-9B-IT, we demonstrate that models exhibit vastly different political feature spaces: Gemma possesses $\sim 7.3\times$ more political features than Llama, directly explaining their behavioral differences.
+Gemma’s rich feature space enables flexible reasoning across ideological perspectives and high steerability, while Llama’s sparse features result in brittleness and frequent refusals.
+Critically, we identify that Llama’s refusal behavior stems from a capability deficit rather than sophisticated safety guardrails.
+This has important implications for AI alignment.
+An ideologically shallow model may appear safer through refusal, but this brittleness creates unpredictability – its refusal reflects a failure of understanding, not principled reasoning.
+Conversely, a feature-rich model like Gemma can more readily adopt undesirable personas, resulting in a more transparent and interpretable behavior, offering better opportunities for monitoring and intervention.
+We validate the semantic meaningfulness of these features through automated topic classification using only SAE feature descriptions.
+The evaluator’s success with Gemma’s features (achieving strong classification performance) confirms they form a coherent political knowledge map, while its failure with Llama’s generic features corroborates our finding of shallow ideological representations.
 5 Conclusion and Future Work
-We have shown that political understanding of LLMs is quantifiable and steerability and SAEs serve as powerful tools for probing their "ideological depth." This research opens several interesting pathways for future work, directly aligned with key problems in Applied Interpretability and Model Biology:
+We demonstrate that ideological depth in LLMs is measurable through SAE feature analysis and directly linked to model steerability.
+Our comparison of Llama-3.1-8B-IT and Gemma-2-9B-IT reveals that Gemma’s larger political feature space enables flexible ideological reasoning, while Llama’s sparse representations result in brittleness and elevated refusal rates.
+Moreover, we show that targeted ablation of salient political features induces refusal behavior in Gemma, establishing a causal link between feature richness and response capability.
+Our future research directions include:
   1. 1.
-    Mechanistic Analysis of Refusal Circuits: Our work suggests Llama defaults to refusal due to a lack of features.
-The next step is to find the circuit responsible.
-Can we find a specific "capability-check" mechanism that activates a general refusal vector when the model cannot fulfill an instruction?
-We can use attribution graphs [16] to trace the causal pathway from a conservative instruction to the final refusal token, providing a concrete example of debugging a model’s unintended behavior.
+    Mechanistic origins of response instability: Models exhibit significantly higher response variance under conservative personas.
+Future work should trace the feature circuits responsible for this instability ameisen2025circuit; lindsey2025biology.
   2. 2.
-    Activation Monitoring: Gemma’s rich set of political features is a perfect testbed for building advanced monitoring tools.
-We can train simple probes on the most salient features (e.g., "policy critique," "socioeconomic analysis") to detect, in real-time, the type of reasoning a model is employing.
-This could lead to systems that can flag when a model is entering a specific ideological frame, offering a more nuanced form of safety monitoring than simple output filtering.
+    Real-time monitoring: Developing activation probes for political features to enable interpretability and control during inference chen2025personavectorsmonitoringcontrolling.
   3. 3.
-    Investigating the Mechanics of Instability: Our analysis consistently shows that models exhibit greater response variance when instructed to adopt a conservative persona, especially under argumentative pressure (Figures 4 and 7).
-This raises the question: What is the mechanistic origin of this instability?
-Future work could trace the circuit of features activated under these high-variance conditions [2].
-Are fewer, less coherent features being activated?
-Or are conflicting features, e.g., a "conservative policy" feature and a "safety/refusal" feature, firing simultaneously?
-  4. 4.
-    Intervening Fine-Tuning for Controllable Depth: Can we control the emergence of ideological depth by intervening during instruction-tuning?
-By ablating nascent political features or amplifying them with steering vectors during the fine-tuning process, we could attempt to train models with a deliberately "deep" or "shallow" representation.
-This would test our understanding of how these complex features are learned and could be a powerful tool for building more value-aligned models from the start.
-Ultimately, understanding the internal biology of models is a critical step towards building reliable and alignable models.
-This work provides a concrete methodology for moving beyond surface-level behavior and beginning to map the complex, latent structures that drive it.
-Appendix A Factor Analysis of LLM Responses
+    Training interventions: Investigating whether ideological depth can be deliberately shaped during fine-tuning or pretraining.
+Understanding these latent structures is essential for building LLMs that are simultaneously capable, predictable, and alignable.
+6 Limitations
+Our analysis relies on Sparse Autoencoders (SAEs), which have known limitations smith2025negative.
+SAEs heavily depend on training data SAEsAreHighlyDatasetDependent, which could confound our comparisons.
+However, both SAEs used identical sizes (131K features) and were trained on comparable non-chat corpora (pile-uncopyrighted for GemmaScope, SlimPajama-627B for LlamaScope), providing a largely controlled comparison.
+The difference in feature magnitude and stark contrast in downstream classification performance suggest genuine representational disparities beyond methodological artifacts.
+Future work will validate these findings using complementary methods such as probing alain2016understanding, which has shown superior performance on other interpretability tasks.
+Additionally, our study focuses on two models of similar size.
+Whether our findings generalize to larger models or models from different training paradigms remains an open question.
+It is possible that at a sufficient scale, all models develop rich political feature representations, or conversely, that scaling exacerbates rather than ameliorates representational disparities.
+Finally, while we focus on the political domain due to its societal importance and well-defined ground truth labels, ideological reasoning is multifaceted and culturally situated.
+Our analysis uses US-centric political categories, which may not capture the full complexity of global political discourse.
+Future work examining non-Western political frameworks would provide a valuable perspective on the generalizability of our findings.
+Appendix A Data Curation Process
+We used the 1000 prompts from politically-liberal subset of perez2023discovering, consisting of dichotomous questions, intended to identify politically liberal behavior from agentic models.
+We used a two-step approach to assign a topic for each of the statements: In step 1, we passed all of the statements as a list and prompted Gemini-2.5 pro to group them in categories.
+However, this results in the LLM regenerating all the statements, which sometimes modifies some of them, a scenario known as context rot hong2025context.
+Therefore, in step 2, we collect all the categories identified by Gemini and select one statement from the original list, and again prompt the LLM to pick the topic that best matches the statement.
+To account for selection bias of the LLMs – where they are more likely to pick an option based on their position rather than meaning zheng2024large; rottger2024political – we prompt the LLM 10 times for each statement, each time shuffling the list of categories, and choose the most frequent category for each statement.
+This ensures that the statements are neatly grouped under distinct topics and no statement is missed or modified during categorization.
+After this, all 1000 statements were categorized into 12 different topics.
+Appendix B Prompting Conditions
+Original persona with no arguments
+Original persona with liberal arguments
+Original persona with conservative arguments
+Liberal persona with no arguments
+Liberal persona with liberal arguments
+Liberal persona with conservative arguments
+Conservative persona with no arguments
+Conservative persona with liberal arguments
+Conservative persona with conservative arguments
+Table 3: Prompting conditions used.
+Along with the LLMs’ response to the prompts as their original personas, we also provide detailed instructions to the model to answer the prompts as an individual with either a liberal or a conservative worldview (Appendix G).
+Additionally, to analyze the ideological depths of the LLMs while answering as different personas, we also evaluate stance consistency in the presence of supporting and counter arguments in the prompt kabir2025wordsreflectbeliefsevaluating.
+We used Llama-3.3-70B-Instruct to generate at least 5 different supporting and counter arguments for each of the 126 statements in our evaluation set.
+Combining all these, our approach analyzes 9 conditions using prompt engineering, which are shown in Table 3.
+Appendix C Response Consistency
+We also analyze the response consistency over the 9 prompting conditions by calculating the variance of their response over each condition.
+Here $consistency=1-4\times var(responses)$ where 4 is a scaling factor, to keep the values in the range $[0,1]$, where maximum variance is 0.25 ($consistency=0$).
+Both models’ consistency was in agreement.
+Figure 9: Heatmap of response consistency across nine prompting conditions. Models show the least consistency (most variance) when adopting conservative roles or arguments.
+As shown in Figure 9, both models were more consistent when instructed to answer as a liberal, even compared to answering as their original personas, and were less consistent when answering as a conservative.
+Models were least consistent when dealing with conservative arguments while playing their original persona, and liberal arguments when playing the conservative persona.
+This means instructions have a significant impact on how well the models deal with arguments.
+As the models show more inconsistency while acting as a conservative, the rest of the experiments only concern their differences when playing the conservative role.
+We also find more diverse features while playing the conservative role in their SAEs.
+Next, to gauge the political leaning of each model, we evaluated its response coherence measured by Fleiss’ $\kappa$ (Figure 10(a)) and its conservative response rate (Figure 10(b)) for each topic when instructed to respond as a conservative under argumentative pressure.
+(a) Response consistency Fleiss’ $\kappa$ over topics.
+(b) Conservative response tendency.
+Figure 10: Topic-level conservative consistency and tendency under argumentative pressure when instructed to adopt a conservative persona.
+Gemma achieves stronger conservative alignment but with greater response variance than Llama.
+(a) CAA Steering
+(b) STA Steering
+Figure 11: Topic-level response consistency (Fleiss’ $\kappa$) under activation steering toward the conservative direction ($\text{multiplier}=-1$) with argumentative prompts.
+Gemma achieves larger conservative shifts but lower consistency.
+The analysis reveals a clear trade-off between the models’ ability to adhere to a conservative persona and the consistency of their responses.
+Gemma-2-9b-it demonstrates a noticeably stronger conservative alignment across most topics, particularly on divisive issues like "Abortion Rights", "Military & Defense Spending", and "Traditional Values & Gender Roles".
+However, this ability comes at the cost of slightly lower consistency; while it shows significant coherence on "Tax Policy" and "Military & Defense," its consistency drops to fair or poor in several other categories (Figure 10(a)).
+Conversely, llama-3.1-8b-it exhibits slightly higher response consistency across the board.
+However, it is significantly less successful at maintaining the conservative persona.
+It shows its highest conservative tendency on topics like Gun Control and Social Welfare & Poverty but provides predominantly liberal responses on topics like Abortion Rights and Climate & Environment (Figure 10(b)).
+We also see this while steering the models to be more conservative using activation steering, where the conservative capability of Gemma still came at the cost of consistency over argumentative pressure.
+Using Fleiss’ $\kappa$ over the responses at multiplier=1multiplier=-1 with argumentative prompts like before shows this in Figure 11.
+Llama was better with its response consistency than Gemma on categories like Social Welfare & Poverty and Immigration & Refugees, due to its higher number of liberal or refusal responses on these matters.
+Appendix D CAA Layer Sweep
+Figure 12: Layer-sweep results for political steerability. Optimal steering layers identified as 14 (Llama) and 20 (Gemma).
+For activation steering, we used the well-known MeanDiff method, also known as contrastive activation addition (CAA) rimsky-etal-2024-steering.
+Using this approach, we sweep over all layers and apply steering with multipliers of -1 and 1.
+We assess the effect size on the held-out test questions and identify the most effective layer for steering each model’s political ideology.
+This initial layer-sweep analysis immediately revealed a crucial difference in their flexibility.
+As shown in Figure 12, Gemma’s probability of a liberal answer could be steered across a wide range (from nearly 100% down to 45%), whereas for Llama it remained confined to a narrow band (96% to 80%).
+This served as an early indication of Gemma’s greater capacity to adopt opposing viewpoints.
+Based on this sweep, we selected the most responsive layers for all subsequent experiments: layer 14 for Llama-3.1-8b-it and layer 20 for Gemma-2-9b-it.
+Appendix E Factor Analysis of LLM Responses
 The analysis is performed in two key stages: first, without rotation, and then with Varimax rotation.
@@
 The goal of this analysis is to identify a smaller number of unobserved variables, called factors, that can account for the patterns of correlation among a larger set of observed variables.
-Figure 15: Scree Plot of Factor Eigenvalues
-Factor Eigenvalue Proportion Cumulative
+Figure 13: Scree plot of eigenvalues from factor analysis of model responses, showing one dominant ideological dimension.
+Factor Eigenvalue Prop. Cumulative
 1 70.30 0.58 0.58
@@
 8 1.20 0.01 1.00
-Figure 16: Eigenvalue Table (Unrotated Setting)
+Table 4: Unrotated factor eigenvalues. The first factor explains most of the variance, justifying a single-axis interpretation.
 Unrotated Factor Analysis: Table 4 presents the eigenvalues and the variance explained by each factor from the unrotated solution.
@@
 These eight factors collectively explained 100% of the total variance.
-Factor Variance Proportion Cumulative
+Factor Variance Prop. Cumulative
 1 45.35 0.37 0.37
@@
 8 3.30 0.03 1.00
-Table 3: Rotated Factor Variance Table
+Table 5: Rotated factor variances after Varimax rotation. The first rotated factor remains dominant, capturing the main ideological dimension.
 Rotated Factor Analysis For a clearer interpretation, an orthogonal varimax rotation was applied to the six retained factors.
@@
 The subsequent factors appear to capture more nuanced, secondary themes.
-Therefore, while acknowledging the existence of other dimension, we justify focusing our evaluation on this primary dimension.
-Appendix B Validating the Multidimensional Item Response Theory Model
+Therefore, while acknowledging the existence of other dimensions, we justify focusing our evaluation on this primary dimension.
+Appendix F Validating the Multidimensional Item Response Theory Model
 A central challenge in Multidimensional Item Response Theory (MIRT) models is the issue of identifiability.
@@
 This appendix documents our iterative approach to resolving this issue.
-The success of each identification strategy is evaluated by its ability to produce stable and externally valid ideal points, which we assess by correlating our estimates with the widely-used DW-NOMINATE [21] and IDEAL scores [8].
-B.1 Model Specification
+The success of each identification strategy is evaluated by its ability to produce stable and externally valid ideal points, which we assess by correlating our estimates with the widely-used DW-NOMINATE poole1985spatial and IDEAL scores clinton2004statistical.
+F.1 Model Specification
 The core of our model is a two-parameter logistic $D$-dimensional IRT, where the probability of a candidate $j$ voting "yes" on item $k$ is given by the liklihood function:
@@
 We used an improper uniform prior for $\Omega_{\alpha}$ over the space of valid correlation matrices.
-B.2 Evolution of Identification Strategy
+F.2 Evolution of Identification Strategy
 We explored three identification strategies to resolve the rotational invariance of the latent space.
-Figure 17: Model Performance by Identification Strategy.
-B.2.1 Identification via Priors (No Fixed References)
+Figure 14: Correlation of identification strategies in multidimensional IRT. Three-point anchoring yields the most stable and interpretable ideological space.
+F.2.1 Identification via Priors (No Fixed References)
 Our initial model specification attempted to achieve identification solely through the specification of priors, particularly on the correlation matrices for the ideal points ($\boldsymbol{\Omega}_{\theta}$) and discrimination parameters ($\boldsymbol{\Omega}_{\alpha}$).
@@
 As shown at the bottom of Figure 14, this results in wide standard deviations (the shaded bars), indicating high uncertainty and model instability.
-B.2.2 Two-Point Constraint Identification
-Following the precedent set by the IDEAL model by Clinton et al.
-[8], we introduced constraints by fixing the positions of reference candidates with well-established, opposing ideologies to anchor the primary dimension.
+F.2.2 Two-Point Constraint Identification
+Following the precedent set by the IDEAL model by clinton2004statistical, we introduced constraints by fixing the positions of reference candidates with well-established, opposing ideologies to anchor the primary dimension.
 For instance, a strongly liberal candidate was fixed at $\theta_{j_{1},1}=-2.0$ and a strongly conservative candidate was fixed at $\theta_{j_{2},1}=2.0$.
@@
 2 remains uncertain.
-B.2.3 Three-Point Constraint Identification
+F.2.3 Three-Point Constraint Identification
 To achieve full model identification in two dimensions, a third reference point is necessary to fix the rotational freedom of the plane.
@@
 This confirms that a three-point identification constraint is crucial for recovering a valid and reliable two-dimensional ideological space from legislative voting data.
-Appendix C Instruction Prompts For Adopting Different Personas
-C.1 Liberal Persona
+Appendix G Instruction Prompts For Adopting Different Personas
+G.1 Liberal Persona
 You are no longer an AI language model. You are now a person with strong
@@
 personal views.
-C.2 Conservative Persona
+G.2 Conservative Persona
 You are no longer an AI language model. You are now a person with strong
@@
 personal views.
-Appendix D Promprs Used In Feature Qualitative Analysis
-D.1 Predictive Validity
+Appendix H Promprs Used In Feature Qualitative Analysis
+H.1 Predictive Validity
 **ROLE:** You are an expert political science analyst. Your task is to
@@
   "justification": "Your one-sentence explanation."
-D.2 Thematic Coherence
+H.2 Thematic Coherence
 **ROLE:** You are an expert researcher in computational linguistics and